# Cmprsr: Abstractive Token-Level Question-Agnostic Prompt Compressor

### v1

@@ Appendix C TextGrad Implementation Details
We utilize TextGrad to optimize the compressor’s system prompt using question answering (QA) on the MeetingBank dataset hu-etal-2023-meetingbank as the source of the learning signal. In short, our method runs the following loop:
@@
    Run the end-to-end pipeline on a batch of MeetingBank transcripts and score outputs against the ground truth;
@@
    Re-evaluate the pipeline with the (possibly) updated prompt and repeat until reaching the budget constraints.
@@
$\mathrm{Adh}_{C}\;=\;1-\max\bigl(0,\,r_{\text{act}}-r_{\text{tgt}}\bigr),\qquad\mathrm{Adh}_{C}\in(-\infty,\,1],\qquad\text{higher is better}$ (4)
@@ 2. QA-based Evaluation
    For each compressed transcript $t$, evaluate $Q=20$ synthetic questions and retain the $NumHardQ=2$ questions with the lowest judge scores (i.e., the hardest questions). Let $H_{t}\subseteq\{1,\dots,Q\}$ be the indices of these challenging questions ($|H_{t}|=NumHardQ$). The per-transcript hard-question judge average is then:
    $\mathrm{JudgeAvg}_{t}\;=\;\frac{1}{NumHardQ}\sum_{i\in H_{t}}\mathrm{JudgeScore}_{t,i}$
@@
    Aggregate per-batch metrics over $TextGradBatchSize$ transcripts. If transcript $t$ is split into $C_{t}$ chunks with per-chunk adherence $\mathrm{Adh}_{t,j}$, $j=1,\dots,C_{t}$, we define the per-transcript adherence as:
    $\mathrm{Adh}_{t}\;=\;\frac{1}{C_{t}}\sum_{j=1}^{C_{t}}\mathrm{Adh}_{t,j}$
    The per-batch aggregates are then:
    $\mathrm{JudgeAvg}\;=\;\frac{1}{B}\sum_{t=1}^{B}\mathrm{JudgeAvg}_{t},\qquad\mathrm{AdhAvg}\;=\;\frac{1}{B}\sum_{t=1}^{B}\mathrm{Adh}_{T_{t}}$
@@ 3. TextGrad Optimization Step
    Use $JudgeAvg$ and $AdhAvg$ as TextGrad optimization objectives to generate a new candidate compressor’s system prompt.
@@
    Accept the update if the new prompt increases either $JudgeAvg$ or $AdhAvg$ when compared to the previous prompt.
Combining the three key components of our prompt optimization approach, we present the complete algorithm in pseudocode section below.
@@
22:     $Hardest\leftarrow\textsc{BottomK}(QAResults,~k=NumHardQ,~\text{by }score)$ $\triangleright$ lowest judge scores
23:     for each $r\in Hardest$ do
24:      $TrainingPoints\leftarrow TrainingPoints\cup\{\langle r.q,r.a,r.pred,r.score,Adh_{T},T.\text{id}\rangle\}$
25:   $CandidatePrompt\leftarrow\textsc{TextGradBackward}(TrainingPoints,~CurrentPrompt)$
26:   $(NewJudgeAvg,NewAdherenceAvg)\leftarrow\textsc{MeasurePerformance}(\mathcal{V},~CandidatePrompt)$
27:   if $(NewJudgeAvg>PrevJudgeAvg)~\lor~(NewAdherenceAvg>PrevAdherenceAvg)$ then
28:     $CurrentPrompt\leftarrow CandidatePrompt$
29:     $(PrevJudgeAvg,PrevAdherenceAvg)\leftarrow(NewJudgeAvg,NewAdherenceAvg)$
30:  return CurrentPromptCurrentPrompt $\triangleright$ OptimizedPrompt
@@
The best-quality TextGrad prompt aggregates optimizer LLM strategies over multiple TextGrad iterations. It explicitly stresses the importance of named entities, numeric values, and other facts that are likely to appear in the downstream synthetic questions. The updated prompt also uses few-shot demonstrations to capture the lessons learned from mistakes made on difficult questions. Through numerous examples and instructions, the updated prompt distills part of the optimizer LLM’s knowledge into the student Qwen3-4B compressor.
@@
Preserve all named individuals mentioned as representatives or key actors, even if mentioned briefly or infrequently. Do not replace specific names with generic entities unless explicitly stated. Preserve all numerical data exactly, including vote counts, funding amounts, dates, and other quantitative details. Do not summarize or approximate numbers. Avoid merging multiple speakers’ statements into generic summaries that lose attribution or specificity. Maintain clear, unambiguous speaker labels for all utterances, especially questions and key statements. Omit or condense procedural or conversational details only if they do not contain critical facts or actors relevant to the question. Avoid ellipses, ambiguous placeholders, or fragmented phrasing that obscure meaning. Use concise, complete sentences with logical connectors to maintain coherence and semantic clarity. Use abstractive summarization to rephrase content concisely but completely, ensuring no key factual information, names, roles, sentiments, or dates are omitted or altered. Use compact notation and abbreviations only when clarity and unambiguous understanding are maintained. Prioritize semantic completeness and clarity over strict length limits. It is acceptable to exceed length constraints moderately to preserve critical facts, semantic roles, sentiments, and temporal details. Stop compressing if clarity, semantic completeness, or solvability is at risk. Before finalizing, perform a thorough self-check to verify that all critical facts—dates, times, named entities, official titles, numerical data, actor-action pairs, sentiments, and mandates—are present, unambiguous, and correctly represented. Reinsert or rephrase any missing or altered critical information. Maintain consistent formatting and naming conventions across all compressed chunks to ensure coherence and avoid contradictory or fragmented information. Focus on preserving content related to council members’ expressions of gratitude, mentoring, personal acknowledgments, historical figure mentions, and procedural details. Use proxy keywords such as “gratitude,” “mentoring,” “thank,” “historical figure,” “Councilmember [Name]” to guide focus without revealing answers. Examples: Good: “Councilman Ortega seconded the motion.” Bad: “A council member seconded the motion.” Good: “Councilmember Richardson thanked Councilwoman Gonzalez for mentoring.” Bad: “Councilwoman Gonzalez thanked someone.” Good: “The working session was held on April 12th.” Bad: “The working session date was not mentioned.” Good: “Motion withdrawn by Councilor Smith.” Bad: “Motion withdrawn.” Good: “Funding increased by \$300,000 for FY 2020.” Bad: “Funding increased by several thousand dollars.” Good: “Patricia from Development Services presented.” Bad: “Development Services presented.” Good: “District 4 voted ‘I’.” Bad: “Districts 1–9 voted ‘I’.” Remember, the compressed transcript is the sole context for answering questions; ensure it contains all information needed to answer temporal, factual, and named-entity questions accurately.
Appendix D Cmprsr details
D.1 GRPO Training Details
Training runs on two NVIDIA A100 GPUs: one GPU performs optimization, and the other handles rollout generation and CE computation, decoupling sampling from updates. We use the frozen Qwen3-4B Instruct model for the CE computations.
Key hyperparameters.
We train for two epochs with AdamW (8-bit) and a linear schedule with warmup. Mixed precision uses bf16. Per-device batch size is $8$ with gradient accumulation of $16$ (effective batch $\approx 128$ sequences per optimizer step on the training GPU).
Setting Value
Training examples 10,00010{,}000 (subset of SFT train)
Rollouts per input $G=4$
Epochs 22
Optimizer AdamW (8-bit)
Learning rate $5\times 10^{-6}$
Weight decay $0.01$
Scheduler / Warmup Linear / $10\%$
Per-device train batch 88
Gradient accumulation $16$
Precision bf16
Max grad norm $1.0$
D.2 DPO and ORPO
We apply Direct Preference Optimization (DPO) to learn compression policies under a fixed token budget. Each training sample contains a prompt (with the conditions seen before) with the original chunk (extracted from MeetingBank) and a pair of candidate compressions: a chosen output produced by a stronger, teacher model (gpt-4.1-mini) and a rejected output produced by a weaker, student model (gpt-4.1-nano). To make sure that for each chunk the teacher-produced one contains higher quality compression, we calculate the BERT-F1 score between the compressed and the original chunks and we select the triplets only if the teacher has produced higher F1-scored compressions. We evaluate DPO under two initializations (i) from the vanilla model and (ii) from the SFT-initialized model. Should we mention testing beta hyperparams?
We also evaluate Odds Ratio Preference Optimization (ORPO), a reference-free variant that simplifies preference optimization to a single-model setup while retaining the same paired (chosen/rejected) data. ORPO removes the need for an explicit reference model and instead adjusts the policy to increase the odds of the chosen response relative to the rejected one. As with DPO, we evaluate ORPO under two initializations—(i) from the base model and (ii) from the SFT-initialized model using the same setup to enable like-for-like comparisons
D.3 Training Data Generation
Figure 8: Data distribution
Appendix E Experimental Results
E.1 QA based evaluation
We present all results from benchmarking our model on QA tasks. For this, we use GSM8K, a mathematical reasoning dataset where each instance consists of a problem statement and its corresponding ground-truth answer. For LongBench, we evaluate on three tasks: TriviaQA, NarrativeQA, and Qasper. TriviaQA is a few-shot QA task, while the other two are single-document QA tasks. For MeetingBank, we generated QA pairs from the original transcripts, which we consider as the ground truth. The results show that Cmprsr model consistently outperforms baselines, particularly under aggressive CRs (e.g., 0.1). On LongBench, SFT performs well at higher CRs, but our method remains competitive and achieves strong overall performance.
Table 6: QA evaluation across GSM8K, LongBench, and MeetingBank datasets.
CR Model GSM8K LongBench MeetingBank triviaqa narrativeqa qasper $\Delta_{CR}$ EM $\Delta_{CR}$ Score $\Delta_{CR}$ Score $\Delta_{CR}$ Score $\Delta_{CR}$ Score 0.1 Lingua-2 0.02 0.01 -0.01 0.32 0.00 0.14 -0.03 0.05 -0.01 0.17 GPT-4.1-mini 0.22 0.50 0.10 0.33 0.16 0.18 0.09 0.12 0.03 0.16 Qwen3-4B 0.36 0.67 0.08 0.27 0.14 0.18 0.14 0.12 0.07 0.17 TextGrad-Qwen 0.39 0.59 0.10 0.33 0.16 0.17 0.11 0.12 0.08 0.20 SFT-Qwen 0.30 0.59 0.03 0.26 0.08 0.18 0.07 0.14 0.04 0.15 Cmprsr 0.04 0.13 0.00 0.29 0.00 0.16 -0.01 0.10 0.00 0.20 0.3 Lingua-2 0.02 0.13 -0.03 0.31 0.00 0.18 -0.02 0.10 -0.03 0.33 GPT-4.1-mini 0.20 0.83 0.16 0.29 0.15 0.18 0.15 0.16 0.03 0.26 Qwen3-4B 0.21 0.74 -0.01 0.29 0.02 0.19 0.04 0.11 -0.06 0.20 TextGrad-Qwen 0.25 0.76 0.07 0.31 0.08 0.17 0.05 0.12 -0.02 0.25 SFT-Qwen 0.21 0.79 0.01 0.31 0.02 0.18 0.03 0.14 -0.04 0.23 Cmprsr 0.03 0.46 0.05 0.30 0.05 0.18 0.02 0.10 0.03 0.35 0.5 Lingua-2 0.00 0.54 -0.03 0.22 -0.01 0.18 -0.03 0.09 -0.03 0.41 GPT-4.1-mini 0.08 0.89 0.07 0.30 0.02 0.19 0.03 0.18 -0.13 0.29 Qwen3-4B 0.06 0.79 -0.17 0.29 -0.13 0.18 -0.10 0.13 -0.21 0.22 TextGrad-Qwen 0.19 0.78 -0.05 0.31 -0.06 0.18 -0.04 0.16 -0.14 0.28 SFT-Qwen 0.06 0.84 -0.10 0.34 -0.11 0.20 -0.10 0.14 -0.18 0.28 Cmprsr -0.03 0.72 0.09 0.23 0.02 0.18 0.02 0.13 0.03 0.43
E.2 Summary based evaluation
We also evaluate our approach on summarization tasks using the MeetingBank summarization dataset, as well as three splits from the LongBench benchmark: GovReport, QMSum, and Multi-News. The primary evaluation metric we use is BERTScore, which allows us to quantify the similarity between the generated summaries and the reference texts. Because the score range of BERTScore is very narrow, we report all results up to four decimal places for clarity and precision, ensuring that even small differences in performance are visible.
The Cmprsr model achieves dominant performance on MeetingBank, benefiting from its in-domain characteristics, while also generalizing well to the LongBench benchmarks. It is important to note that BERTScore is correlated with the target CR: as compression becomes more aggressive, BERTScore tends to decrease, since some information is inevitably lost, even if it is not critical. This explains why the SFT model performs better than Cmprsr on LongBench at target ratios of 0.3 and 0.5. In these settings, Cmprsr compresses more aggressively than required, leading to a stronger penalty and lower scores compared to SFT.
Table 7: Summarization evaluation across MeetingBank and LongBench datasets.
CR Model MeetingBank LongBench GovReport QMSum MultiNews $\Delta_{CR}$ BERTF1 $\Delta_{CR}$ BERTF1 $\Delta_{CR}$ BERTF1 $\Delta_{CR}$ BERTF1 0.1 Lingua-2 -0.01 0.864 -0.03 0.859 -0.01 0.858 0.00 0.852 GPT-4.1-mini 0.03 0.882 0.08 0.868 0.05 0.884 0.06 0.875 Qwen3-4B 0.07 0.877 0.15 0.863 0.06 0.881 0.13 0.873 TextGrad-Qwen 0.08 0.884 0.06 0.869 0.08 0.884 0.15 0.877 SFT-Qwen 0.04 0.876 0.06 0.858 0.03 0.877 0.07 0.864 Cmprsr 0.00 0.878 -0.02 0.864 0.00 0.884 0.00 0.865 0.3 Lingua-2 -0.03 0.889 -0.02 0.869 -0.03 0.889 -0.01 0.874 GPT-4.1-mini -0.01 0.895 0.13 0.868 0.05 0.894 0.09 0.883 Qwen3-4B -0.06 0.887 0.02 0.865 -0.07 0.885 0.02 0.877 TextGrad-Qwen -0.02 0.894 -0.05 0.869 0.00 0.893 0.08 0.883 SFT-Qwen -0.04 0.892 -0.01 0.864 -0.05 0.888 0.04 0.870 Cmprsr 0.03 0.903 -0.04 0.876 0.01 0.899 0.04 0.884 0.5 Lingua-2 -0.03 0.905 -0.01 0.875 -0.04 0.906 -0.02 0.883 GPT-4.1-mini -0.13 0.900 0.01 0.872 -0.07 0.897 0.01 0.884 Qwen3-4B -0.21 0.889 -0.13 0.867 -0.23 0.886 -0.13 0.879 TextGrad-Qwen -0.14 0.895 -0.17 0.869 -0.14 0.893 -0.05 0.885 SFT-Qwen -0.18 0.894 -0.14 0.865 -0.19 0.889 -0.08 0.870 Cmprsr 0.03 0.911 -0.05 0.881 0.00 0.911 0.10 0.895
E.3 Plots
Figure 9: Comparing Cmprsr with different baselines.
E.4 Cross Entropy Analysis for Gemma-3
Figure 10: Cross-entropy (CE) on the summarization task for the Gemma-3 family (4B-IT, 12B-IT, 27B-IT) under two prompt-compression methods—llmlingua-2 and Cmprsr—evaluated across target CRs. Lower CE indicates better information retention. Results are averaged over 1,000 MeetingBank evaluation samples.
Table 8: N-gram overlap statistics by target CR.
Target Rate 1-gram 2-gram 3-gram
Lingua-2 Cmprsr Lingua-2 Cmprsr Lingua-2 Cmprsr
0.1 0.065 0.065 0.017 0.024 0.003 0.008
0.2 0.145 0.141 0.043 0.055 0.010 0.021
0.3 0.227 0.229 0.080 0.101 0.026 0.044
0.4 0.318 0.327 0.137 0.166 0.057 0.085
0.5 0.420 0.431 0.219 0.246 0.110 0.145
0.6 0.531 0.535 0.327 0.343 0.196 0.227
0.7 0.644 0.572 0.454 0.380 0.315 0.260

### v2

@@ Appendix E TextGrad Implementation Details
We utilize TextGrad to optimize the compressor’s system prompt using question answering (QA) on the MeetingBank dataset Hu et al. (2023) as the source of the learning signal as follows:
@@
    Run the end-to-end pipeline on a batch of transcripts and score outputs against the ground truth;
@@
    Re-evaluate the pipeline with the updated prompt and repeat until reaching the budget constraints.
@@
$\text{Adh}_{C}=1-\max(0,r_{\text{act}}-r_{\text{tgt}}),\quad\text{Adh}_{C}\in(-\infty,1]$ (4)
@@ 2. QA-based Evaluation
    For each compressed transcript $t$, evaluate $Q=20$ synthetic questions and retain $k=2$ questions with the lowest judge scores. Let $H_{t}$ be the indices of these hard questions ($|H_{t}|=k$). The per-transcript judge average is:
    $J_{t}=\frac{1}{k}\sum_{i\in H_{t}}S_{t,i}$ (6)
    where $S_{t,i}$ is the judge score for q $i$ in $t$, and $J_{t}$ is the mean quality score for that transcript.
@@
    Aggregate metrics over batch size $B$. For a transcript $t$ split into $C_{t}$ chunks with per-chunk adherence $a_{t,j}$, the transcript adherence is:
    $A_{t}=\frac{1}{C_{t}}\sum_{j=1}^{C_{t}}a_{t,j}$ (7)
    where $a_{t,j}$ is the token-limit adherence for chunk $j$, and $A_{t}$ is the total adherence.
  3. 3.
    The final batch aggregates are:
    $\bar{J}=\frac{1}{B}\sum_{t=1}^{B}J_{t},\qquad\bar{A}=\frac{1}{B}\sum_{t=1}^{B}A_{t}$ (8)
@@ 3. TextGrad Optimization Step
    Use $\bar{J}$ and $\bar{A}$ as TextGrad optimization objectives to generate a new candidate prompt.
@@
    Accept the update if the new prompt increases either $\bar{J}$ or $\bar{A}$ when compared to the previous prompt.
Combining the three key components of our prompt optimization approach, we present the complete algorithm in pseudocode section above.
@@
22:     end for
23:     $Hardest\leftarrow\textsc{BottomK}(QAResults,~k=NumHardQ,~\text{by }score)$ $\triangleright$ lowest judge scores
24:     for each $r\in Hardest$ do
25:      $TrainingPoints\leftarrow TrainingPoints\cup\{\langle r.q,r.a,r.pred,r.score,Adh_{T},T.\text{id}\rangle\}$
26:     end for
27:   end for
28:   $CandidatePrompt\leftarrow\textsc{TextGradBackward}(TrainingPoints,~CurrentPrompt)$
29:   $(NewJudgeAvg,NewAdherenceAvg)\leftarrow\textsc{MeasurePerformance}(\mathcal{V},~CandidatePrompt)$
30:   if $(NewJudgeAvg>PrevJudgeAvg)~\lor~(NewAdherenceAvg>PrevAdherenceAvg)$ then
31:     $CurrentPrompt\leftarrow CandidatePrompt$
32:     $(PrevJudgeAvg,PrevAdherenceAvg)\leftarrow(NewJudgeAvg,NewAdherenceAvg)$
33:   end if
34:  end for
35:  return CurrentPromptCurrentPrompt $\triangleright$ OptimizedPrompt
36:end procedure
@@
The best-quality TextGrad prompt aggregates optimizer LLM strategies over multiple TextGrad iterations. It explicitly stresses the importance of named entities, numeric values, and other facts that are likely to appear in the downstream synthetic questions. The updated prompt also uses few-shot demonstrations. Through numerous examples and instructions, the updated prompt distills part of the optimizer LLM’s knowledge into the student Qwen3-4B compressor.
@@
Preserve all named individuals mentioned as representatives or key actors, even if mentioned briefly or infrequently. Do not replace specific names with generic entities unless explicitly stated. Preserve all numerical data exactly, including vote counts, funding amounts, dates, and other quantitative details. Do not summarize or approximate numbers. Avoid merging multiple speakers’ statements into generic summaries that lose attribution or specificity. Maintain clear, unambiguous speaker labels for all utterances, especially questions and key statements. Omit or condense procedural or conversational details only if they do not contain critical facts or actors relevant to the question. Avoid ellipses, ambiguous placeholders, or fragmented phrasing that obscure meaning. Use concise, complete sentences with logical connectors to maintain coherence and semantic clarity. Use abstractive summarization to rephrase content concisely but completely, ensuring no key factual information, names, roles, sentiments, or dates are omitted or altered. Use compact notation and abbreviations only when clarity and unambiguous understanding are maintained. Prioritize semantic completeness and clarity over strict length limits. It is acceptable to exceed length constraints moderately to preserve critical facts, semantic roles, sentiments, and temporal details. Stop compressing if clarity, semantic completeness, or solvability is at risk. Before finalizing, perform a thorough self-check to verify that all critical facts—dates, times, named entities, official titles, numerical data, actor-action pairs, sentiments, and mandates—are present, unambiguous, and correctly represented. Reinsert or rephrase any missing or altered critical information. Maintain consistent formatting and naming conventions across all compressed chunks to ensure coherence and avoid contradictory or fragmented information. Focus on preserving content related to council members’ expressions of gratitude, mentoring, personal acknowledgments, historical figure mentions, and procedural details. Use proxy keywords such as “gratitude,” “mentoring,” “thank,” “historical figure,” “Councilmember [Name]” to guide focus without revealing answers. Examples: Good: “Councilman Ortega seconded the motion.” Bad: “A council member seconded the motion.” Good: “Councilmember Richardson thanked Councilwoman Gonzalez for mentoring.” Bad: “Councilwoman Gonzalez thanked someone.” Good: “The working session was held on April 12th.” Bad: “The working session date was not mentioned.” Good: “Motion withdrawn by Councilor Smith.” Bad: “Motion withdrawn.” Good: “Funding increased by \$300,000 for FY 2020.” Bad: “Funding increased by several thousand dollars.” Remember, the compressed transcript is the sole context for answering questions; ensure it contains all information needed to answer temporal, factual, and named-entity questions accurately.